OpenStudy (anonymous):

Find all solutions of y''+(2/x)y'=1 ; x>0

OpenStudy (anonymous):

ok. so first step simplify.

OpenStudy (anonymous):

see ya.

OpenStudy (anonymous):

@phi @amistre64 @experimentX

OpenStudy (phi):

Any hints on how to solve it ?

OpenStudy (anonymous):

No, just the problem

OpenStudy (anonymous):

I got it! On a side note, do you know how to solve complex roots of characteristic equations such as: (t^2)y''-(4t)y'+6y=0 using Euler equations?

OpenStudy (phi):

What is the solution?

OpenStudy (phi):

or, the approach ?

OpenStudy (anonymous):

I looked at how Wolfram alpha does it. They substitute v=y' and v'=y'' then is just becomes a typical equation, but the answer it gives has two constants and I wasn't sure if that was correct

OpenStudy (phi):

yes, it (finally) came to me. find the integrating factor in the usual way, it is x^2 x^2 y'' + 2x y' = x^2 we know d/dx (x^2 y') = x^2 d/dx y' + y' d/dx(x^2) = x^2 y'' + y' 2x the left side is an exact differential: d (x^2 y') = x^2 dx integrate both sides x^2 y' = x^3/3 + C y' = x/3 + C x^-2 integrate again: y = x^2/6 + - C/x + D C is an unknown constant, so we could just as well write y= x^2/6 + C/x + D